Solve by factoring and using the principle of zero products. Show all work necessary.
4x^2 - 12x = 16

Answer by richwmiller(17219)

4x^2-12x-16=0
divide by 4
x^2-3x-4=0
since the constant is negative one of the factors is positive
their product is -4 and their sum is -3
[-1 and +4] or [+1 and -4]
but which is negative?
it is -4 since we have to add +1 to get -3x
[+4 and -1 would give us +3x for the middle term]
(x-4)*(x+1)=0
x=4 x=-1